What percent of the area under the normal curve is between z=2.25 and





z=0.58

1
Expert's answer
2022-03-14T17:42:47-0400

"P(0.58<Z<2.25)=P(Z<2.25)-P(Z<0.58)="

"=0.98778-0.71904=0.26874=26.874\%" (from z-table).